what mass of sugar is needed to combine with 250 grams of water to prepare a sugar water mixture that is 10.0% Sugar by mass please explain thoroughly step by step
Answer #1

10% sugar by mass means , 10g sugar is present in 100g mixture .

Means in 100g mixuture , 10g sugar and 90g water .

So , with 90 g water combines = 10g of sugar

With 250g of water will combine = (10/90) × 250 = 27.78g of sugar .

So 27.78 g sugar need to be added with 250g of water to make 10% sugar by mass mixture.  

As after this total misxture = 250+27.78 = 277.78 g

And % of sugar = (27.78/277.78)×100 = 10 %